OpenStudy (anonymous):

Given tan A = .78 find angle A in degrees. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

is it 67.43°?

OpenStudy (anonymous):

that is not what wolfram gets http://www.wolframalpha.com/input/?i=arctan%28.78%29

OpenStudy (anonymous):

Given cos B = .12 find angle B in radians. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

arctan(.78)=.6624

OpenStudy (anonymous):

\[\cos^{-1}(.12)\] and a calculator http://www.wolframalpha.com/input/?i=arccos%28.12%29

OpenStudy (anonymous):

wolfram gives answers both in radians (first one) and degrees (second one)

OpenStudy (anonymous):

Given sinB = 11/18 find angle B in degrees. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

37.67°?

OpenStudy (anonymous):

\[\sin ^{-1}\](11/18)= 37.67 degrees

OpenStudy (anonymous):

yes 37.67 degrees